OpenStudy (anonymous):

Jules kicks a soccer ball off the ground and into the air with an initial velocity of 25 feet per second. Assume the starting height of the ball is 0 feet. Approximately, how long does it take until the soccer ball hits the ground again? How do I set up the equation,i dont want answer...

OpenStudy (anonymous):

i think this will be like a projectile motion but you need the initial direction of the velocity.

OpenStudy (anonymous):

thats all the info i have...

OpenStudy (anonymous):

assumung the ball is hit in vertical direction use the equations for constant acceleration displacement = 0, acceleration = -32 ft /s/s, u = 25 ft/s, t = ? equn is 0 = 25t - 1/2 * 32 * t^2

OpenStudy (anonymous):

im using the equation h(t)=-16t^2+vt+s

OpenStudy (anonymous):

yes - thas ok

OpenStudy (anonymous):

so what do i plug in where for my equation. I tried -16t^2+25t then factored (-4t+5) (4t+5) . That right?

OpenStudy (anonymous):

i know the answer is .8 but cant get that the way ive set it up

OpenStudy (anonymous):

so im doing sumptin wrong

OpenStudy (anonymous):

0 = 25t - 1/2 * 32 * t^2 25t - 16t^2 = o t(25-16t) = 0 t = 25/16 secs

OpenStudy (anonymous):

or 1.56 secs

OpenStudy (anonymous):

is used the equation s = ut + (1/2) a t^2

OpenStudy (anonymous):

0.8 is not correct answer there are 4 constant acceleration equations worth remembering v = u + at s = ut + 0.5at^2 v^2 = u^2 + 2as and s = (u+v)t / 2 whre u = initial velocity v = final velocity s = displacement t = time a = acceleration

OpenStudy (anonymous):

thank you so much. I undertand it now...